OpenStudy (anonymous):

what is the absolute value of -6/24

myininaya (myininaya):

what is the distance from -6/24 to 0?

OpenStudy (anonymous):

i'll say 6/24

jimthompson5910 (jim_thompson5910):

\[\large |-\frac{6}{24}|=\frac{6}{24}=\frac{1}{4}\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

6/24